Let X be a random variable with pdf
f( x) = theta e ^(- theta x); theta > 0, x greater than equal to 0.
Find the moment generating function of X,
and hence find first three moments about
origin.
1
Expert's answer
2020-09-09T19:38:50-0400

Solution


"M(t) = E(exp(tx))"

"=\int exp(tx) \theta exp(-\theta x) dx"

"=\theta\int exp(-x(\theta - t)) dx"

Integrating the above function from 0 to infinity yields:


"\theta /( \theta - t)"

Which is the mgf of f(x)

Answer: M(t)=(θ)/(θ−t)



1st moment is obtained by differentiating M(t) with respect to t and setting t to 0:

By division rule, we obtain:


"dM(t) /(dt) = \theta / (\theta - t)^2"

Setting t=0 yields:

"\theta / (\theta)^2 = 1/ \theta"

Which is the 1st moment



2nd moment is obtained by finding the second derivative of the mgf with respect to t and setting t to 0:

By division rule, we obtain:


"(dM(t))^2/ d^2t = 2 \theta / (\theta - t)^3"

as the second derivative.

Setting t to 0 yields:


"2\theta / (\theta)^3 = 2/ \theta ^2"


Which is the second moment



The 3rd moment is similarly obtained by getting the third derivative of the mgf with respect to t and setting t to 0:

By division rule, we obtain


"(dM(t)) ^3/ dt^3 = 6\theta /(\theta-t) ^4"

as the third derivative.

Setting t to 0 yields :


"6\theta / \theta ^4 = 6 /\theta^3"

Which is the third moment of f(x)